# Inferring fine-grained migration patterns across the United States

### v1

@@ D.2 Validating the estimation procedure using held-out data
Table 4: Validation of estimates in held-out data settings. The full model validations are repeated for comparison. For in-migration rates, metrics are weighted by populations to account for disparate county and state sizes.
We verify that our estimation procedure yields highly correlated estimates with each Census data source even when it is removed from the datasets used for estimation. For example, we remove CBG-level Census populations from our estimation datasets, and verify that our resulting estimates remain highly correlated with CBG-level data. Specifically, we conduct our full battery of validations on two additional sets of migration estimates: “CBG held-out”, which is estimated without the CBG-level population data (i.e. $E^{(t)}$ was not scaled to match ACS 5-year CBG populations); and “State held-out”, which is estimated without the state-level flow data (i.e. $E^{(t)}$ was not scaled to match ACS 1-year state-to-state flows nor state mobility rates). Table 4 reports values for the Pearson correlation across all estimates and ground-truth as well as relative error reduction from Infutor. Values for the original MIGRATE estimates, reported in Figure 2 and Figure 6, are reproduced for easy comparison.
@@
D.3 Related work documenting biases in the Infutor data
@@
As we report in Table 2, Infutor undercounts the national US population. This is consistent with prior work; a similar finding is reported by Qian and Tan [19], Bernstein et al. [17], and Diamond et al. [15] among others. Phillips [16] additionally reports that, relative to 2010 ACS data, Infutor undercounts all state populations, but not by much, and out-of-state movers. A contrasting finding, reported by Diamond et al. [4], is that Infutor might overcount the population according to the 2000 Census in San Francisco. Such a finding might be driven by specific spatiotemporal idiosyncrasies of the San Francisco data in 2000, which is not included in the subset we analyze; however, the bulk of the literature, and all the literature which analyzes the same subset of the data we do, is consistent with our finding that Infutor undercounts the data. Diamond et al. [4] also suggest that Infutor might overcount populations due to lacking information about deaths; this is consistent with our finding that Infutor overcounts the older populations in Figure 3c.
@@
D.4 Assessment of additional demographic biases
@@
Figure 7: Additional analysis of socioeconomic and demographic biases at CBG population level. Variables are taken from the 5-year ACS CBG estimates.
@@
We documented biases in the Infutor dataset geographically (Figure 3a) and on the dimensions of urbanization, age, race, household tenure, and poverty (Figure 3c). Here, we provide additional analyses of demographic bias by sex, educational attainment, and number of children per household (Figure 7). We find that Infutor displays negligible bias on the basis of sex. However, Infutor underrepresents populations with less formal education while overrepresenting populations with more formal education. Populations living in households with no children are overrepresented while those in households with children are underrepresented. This finding is consistent with the Infutor data documentation, which describes a lack of records for people under the age of 18 (therefore, children will be undercounted). MIGRATE essentially eliminates biases along all these demographic dimensions.
Appendix E Case Study: Flows from New York City and Income
@@
Figure 8: Average median income of destination CBGs and counties for New York City out-movers in the 2010-19 period. The large map uses MIGRATE estimates and 5-year ACS CBG data. The inset map shows the corresponding distribution using only publicly available migration data from 5-year ACS county-to-county estimates.
@@
We illustrate the value of MIGRATE estimates in an additional case study: contrasting MIGRATE with publicly available county-to-county migration data when analyzing migration stratified by a demographic or socioeconomic feature. Figure 8 plots the average median income of the CBG to which a person moves according to their previous CBG of residence. To compute this quantity for a given CBG $i$, we average the median income of all destination CBGs $j$ for out-movers of CBG $i$ weighted by the number of out-movers from $i$ to $j$. We perform an analogous computation for the county plot in the inset.
@@
Comparing these results to the data available at county level reveals striking within-county variation. For example, the median income of the county to which people from Manhattan moved was \$$\$60,502$—lower than that of the other four counties in New York City. MIGRATE reveals, however, that movers from some CBGs in Manhattan moved to areas with the highest median income among all New York City CBGs — above \$$\$100,000$ — while movers from other Manhattan CBGs moved to areas with some of the lowest median incomes. Intra-county socioeconomic differences factor heavily in out-migration: Manhattan out-movers with the highest-income destinations tended to reside in the central and lower parts of the county (which tend to have higher incomes), whereas those with lower-income destinations came mostly from northern CBGs (which tend to have lower incomes). Similar within-county dynamics happen in the other four counties. Overall, this analysis illustrates the important demographic heterogeneity in migration which is concealed at the county level.
Appendix F Privacy Protections

### v2

@@ A.2 Validating the estimation procedure using held-out data
Table S1: Validation of estimates in held-out data settings. The full model validations are repeated for comparison. For in-migration rates, metrics are weighted by populations to account for disparate county and state sizes.
We verify that our estimation procedure yields highly correlated estimates with each Census data source even when it is removed from the datasets used for estimation. For example, we remove CBG-level Census populations from our estimation datasets, and verify that our resulting estimates remain highly correlated with CBG-level data. Specifically, we conduct our full battery of validations on two additional sets of migration estimates: “CBG held-out”, which is estimated without the CBG-level population data (i.e. $E^{(t)}$ was not scaled to match ACS 5-year CBG populations); and “State held-out”, which is estimated without the state-level flow data (i.e. $E^{(t)}$ was not scaled to match ACS 1-year state-to-state flows nor state mobility rates). Table S1 reports values for the Pearson correlation across all estimates and ground-truth as well as relative error reduction from Infutor. Values for the original MIGRATE estimates, reported in Figure 2 and Figure S1, are reproduced for easy comparison.
@@
A.3 Related work documenting biases in the Infutor data
@@
As we report in Table M1, Infutor undercounts the national US population. This is consistent with prior work; a similar finding is reported by Qian and Tan [21], Bernstein et al. [19], and Diamond et al. [17] among others. Phillips [18] additionally reports that, relative to 2010 ACS data, Infutor undercounts all state populations, but not by much, and out-of-state movers. A contrasting finding, reported by Diamond et al. [4], is that Infutor might overcount the population according to the 2000 Census in San Francisco. Such a finding might be driven by specific spatiotemporal idiosyncrasies of the San Francisco data in 2000, which is not included in the subset we analyze; however, the bulk of the literature, and all the literature which analyzes the same subset of the data we do, is consistent with our finding that Infutor undercounts the data. Diamond et al. [4] also suggest that Infutor might overcount populations due to lacking information about deaths; this is consistent with our finding that Infutor overcounts the older populations in Figure 3c.
@@
A.4 Assessment of additional demographic biases
@@
Fig. S2: Additional analysis of socioeconomic and demographic biases at CBG population level. Variables are taken from the 5-year ACS CBG estimates.
@@
We documented biases in the Infutor dataset geographically (Figure 3a) and on the dimensions of urbanization, age, race, household tenure, and poverty (Figure 3c). Here, we provide additional analyses of demographic bias by sex, educational attainment, and number of children per household (Figure S2). We find that Infutor displays negligible bias on the basis of sex. However, Infutor underrepresents populations with less formal education while overrepresenting populations with more formal education. Populations living in households with no children are overrepresented while those in households with children are underrepresented. This finding is consistent with the Infutor data documentation, which describes a lack of records for people under the age of 18 (therefore, children will be undercounted). MIGRATE essentially eliminates biases along all these demographic dimensions.
Appendix B Additional analyses of national migration
In this section we provide additional analyses of national migration trends discussed in Section 2.3.
B.1 Out-migration rates by CBG type
Out-migration rate (per 1,000) Out-movers
Plurality Race White 132 31,425,400
Black 128 3,823,677
Asian 124 889,755
Hispanic 118 5,289,710
Urbanization Urban 135 35,481,142
Rural 104 5,947,401
Income Quartile Bottom 130 8,621,155
2nd 130 9,944,217
3rd 131 11,011,339
Top 128 11,851,831
Table S2: Out-migration by group. Out-migration rates (per 1,000 people) and raw counts of out-movers categorized by plurality race, urbanization level, and income quartile. We average the values across all years in the period.
Table S2 reports the average out-migration rates for each of the ten CBG groups we analyze in the main text. The average out-migration rate in all CBGs is $13.0\%$, with $0.9\%$ standard deviation across years. Out-migration rates are lower in plurality Hispanic and rural CBGs; out-migration rates remain similar across income quartiles.
B.2 Additional statistics on migration homophily
Fig. S3: Additional national migration statistics. (a) Flows between the ten types of CBGs discussed in the main text relative to the share of all movers who moved to CBGs in the particular group. Each cell reports the ratio share of movers from origin CBG group moving to destination CBG groupshare of all movers moving to destination CBG $\frac{\text{share of movers from origin CBG group moving to destination CBG group}}{\text{share of all movers moving to destination CBG group}}$; for example, the top left value of $\times1$ reports the ratio of the share of movers from plurality white CBGs who move to plurality white CBGs (90%) to the share of movers from all CBGs who move to plurality white CBGs (80%). (b) Flows between the ten types of CBGs discussed in the main text when restricting to out-of-county movers demonstrate that homophily persists for long-distance moves.
In Figure S3, we extend the analysis of national homophily patterns in migration documented in Figure 4a. Figure S3a shows the shares of movers moving between groups of CBGs relative to the share of all movers who move to that group of CBGs (that is, a version of Figure 4a where the value in each cell is divided by the corresponding value in the row ‘Movers from all CBGs’).
Homophily in migration might occur because many moves are local (within county or within 5 miles) and demographics are spatially correlated. To investigate this possibility, we restrict our analysis to out-of-county moves in Figure S3b. Absolute racial homophily is less prominent—i.e. there are smaller shares in the diagonal of the race by race submatrix—yet still present: out-of-county movers from plurality Black, Asian, and Hispanic CBGs are 4.3$\times$, 5.5$\times$, and 3.1$\times$ likelier than out-of-county movers as a whole to move to CBGs with the same plurality race group (vs. 5.5$\times$, 14.6$\times$, and 4.3$\times$ for all movers).
B.3 Additional statistics on economic mobility in migration
Fig. S4: Further national migration statistics on upwards mobility. Probability of moving to a (a) top-income quartile CBG, (b) bottom-income quartile CBG, and (c) higher-income CBG conditioned on current residence median income and plurality race. In (c), we classify current residence CBGs by income percentiles as opposed to deciles (as in the main text). Disparities remain robust. (d) Movers remain more likely to move to top-income-quartile CBGs than bottom-income-quartile CBGs throughout the decade, with the disparity increasing in magnitude over time.
Figure S4 provides additional analyses supporting our finding of racial disparities in economic mobility (Figure 4b): movers from plurality Asian CBGs are more likely than movers as a whole, and movers from plurality Black CBGs less likely, to move to higher-income CBGs, even when controlling for income of origin CBG. We show that analogous racial disparities still emerge when looking at probability of moving to a CBG in the top income quartile (Figure S4a) or bottom income quartile (Figure S4b) as opposed to the probability of moving to a higher-income CBG (as in the main text). These disparities also emerge when controlling for origin CBG income percentile (Figure S4c), as opposed to decile, as in the main text. We also observe that the gap in upward mobility widens during the decade: the share of movers going to a top income quartile CBG increases from 31% to 33% while the share of movers going to a bottom income quartile CBG decreases from 19% to 17%; overall, the ratio of out-movers moving to top income quartile versus bottom income quartile CBGs increases from 1.64 to 1.95.
B.4 Additional statistics on moving distance
Fig. S5: Further national migration statistics on distance moved. (a) Moving distance by year. (b) Fraction of movers moving within tract, county, and state.
Figure S5 provides additional statistics on moving distance, expanding the findings in Figure 4c. Figure S5a reports how moving distance changes over time, showing a general increase over the decade: the share of movers moving over 50 miles, for example, grows from about 22% to about 26%, and the average moving distance increases from 154 to 191 miles in the decade. Figure S5b reports moving distance stratified by geographic boundary (that is, the share of movers remaining within tract, county, and state). Findings are generally consistent with those reported in Figure 4c, and also with published ACS figures, which state that over half of moves happen within counties [15].
Appendix C Case Study: destinations of New York City outmovers
@@
Fig. S6: Average income of destination CBGs and counties for New York City out-movers in the 2010-19 period. The large map uses MIGRATE estimates and 5-year ACS CBG data. The inset map shows the corresponding distribution using only publicly available migration data from 5-year ACS county-to-county estimates.
@@
We provide a case study contrasting MIGRATE with publicly available county-to-county migration data when analyzing migration stratified by a demographic or socioeconomic feature. Figure S6 plots the average income of the CBG to which a person moves according to their previous CBG of residence. To compute this quantity for a given CBG $i$, we average the mean income of all destination CBGs $j$ for out-movers of CBG $i$ weighted by the number of out-movers from $i$ to $j$. We perform an analogous computation for the county plot in the inset.
@@
Comparing these results to the data available at county level reveals striking within-county variation. For example, the average income of the county to which people from Manhattan moved was \$$\$82,934$—lower than that of the other four counties in New York City. MIGRATE reveals, however, that movers from some CBGs in Manhattan moved to areas with the highest average income among all New York City CBGs — above \$$\$175,000$ at times — while movers from other Manhattan CBGs moved to areas with some of the lowest average incomes. Intra-county socioeconomic differences factor heavily in out-migration: Manhattan out-movers with the highest-income destinations tended to reside in the central and lower parts of the county (which tend to have higher incomes), whereas those with lower-income destinations came mostly from northern CBGs (which tend to have lower incomes). Similar within-county dynamics occur in the other four counties. Overall, this analysis illustrates the important demographic heterogeneity in migration which is concealed at the county level.
Appendix D Case Study: New York City Housing Authority
Ensuring that residents have access to affordable housing is an important and challenging policy goal [68]. One way of achieving this is through the creation and management of public housing units: for example, the New York City Housing Authority (NYCHA) is the largest public housing authority in the US, providing affordable housing to over 500,000 authorized residents [69]. We provide a case study illustrating how the MIGRATE dataset can be used to understand which residents make use of public housing units, thereby illuminating the effect of this component of affordable housing policy.
Fig. S7: We use MIGRATE to understand migration to and from New York City Housing Authority (NYCHA) developments. (a) A map of the 114 Census Block Groups (marked in red) where all residential units are owned by NYCHA (i.e., where all the population lives in a NYCHA development). (b) Time series of in- and out-migration rates for NYCHA CBGs (red line), as well as for CBGs nearby (within 250 meters of a NYCHA development; green line) and non-NYCHA CBGs within New York City (gray line). NYCHA CBGs have consistently lower in- and out-migration rates, implying a less mobile population. (c) Demographics of previous residence CBG for in-movers into NYCHA buildings as well as for the other CBG groups considered, aggregated across the decade (error bars represent standard deviation across years). (d) Demographics of next residence CBG for out-movers.
We identify all 114 New York City Census Block Groups whose residential units are entirely owned by NYCHA using New York City’s Primary Land Use Tax Lot Output (PLUTO) dataset [70] (Fig. S7a). We compare migration patterns for these CBGs, which we refer to as “NYCHA CBGs”, to migration patterns of two other groups of non-NYCHA CBGs: (i) nearby CBGs (within 250 meters of a NYCHA development) and (ii) all non-NYCHA CBGs in New York City. [^10] We excluded from the analysis all 345 CBGs (about 5.3% of all New York City CBGs) where some but not all of the residential units are owned by NCYHA. NYCHA CBGs have had consistently lower in-migration and out-migration rates than the two comparison groups of CBGs throughout the decade (Fig. S7b), implying a less mobile population. This is consistent with prior findings that residents of public housing and recipients of rental benefits have longer tenancies than average New York City residents [69, 71].
To get a sense of the residents being served by NYCHA and their migration trajectories, we examine the demographics of the CBGs where NYCHA in-movers move from (Fig. S7c) and where NYCHA out-movers move to (Fig. S7d). Compared to non-NYCHA in-movers, NYCHA in-movers are likelier to move from plurality Black or Hispanic CBGs, as well as bottom income quartile CBGs. The same is true of NYCHA out-movers: they are also more likely to move to plurality Black or Hispanic CBGs, and bottom-income-quartile CBGs, than are non-NYCHA out-movers. The CBGs that NYCHA movers move to are higher-income than the CBGs they move from, providing evidence of upward mobility. NYCHA out-movers are also more upwardly mobile than non-NYCHA movers: 82% of NYCHA out-movers move to CBGs of higher median income, compared to 55% of all non-NYCHA out-movers.
We also quantify whether NYCHA in-movers and out-movers move within NYCHA, within New York City, and within New York State. There is a significant amount of shuffling within the NYCHA system: 14% of in-movers to a NYCHA CBG came from another NYCHA CBG, and 13% of out-movers from a NYCHA building go to another NYCHA CBG. This shuffling may be due in part to mandated renovations or building closures that result in temporary resident relocations by NYCHA [72, 73, 74]. On average across the decade that we study, NYCHA out-movers are likelier than non-NYCHA movers to remain within New York City (74% vs. 65%); however, this gap has decreased in recent years.
Overall, this analysis highlights how MIGRATE can provide insights into which residents are served by public housing policy. Beyond this case study of New York City, this approach could be extended to study public housing in other cities, or to study the effects of public housing foreclosures—which has been a topic of interest among other migration researchers [18].
Appendix E Privacy Protections